### v1

# Phase driven exact and unconventional superradiance phase transition in non-Hermitian cascaded Rabi cavities

Abstract
This work reports the phase driven symmetry breaking and exact and unconventional superradiance phase transition in the non-Hermitian cascaded Rabi cavities. The non-Hermiticity is introduced in the coupling phase (denoted by $\varphi$) between the atom and the optical field. The exactness refers to the fact that the superradiance phase boundary is obtained analytically and verified by the observables. The unconventionality is reflected in that when $|\varphi|=\frac{\pi}{4}$ or $|\varphi|=\frac{3\pi}{4}$, the phase boundary is uniquely determined by $\mathcal{J}=\frac{1}{2}$ (where $\mathcal{J}$ is the dimensionless cavity coupling strength) and is independent of the atom-optical field coupling strength $g$. For other $\varphi$, the phase boundary is determined by $\mathcal{J}$ and the dimensionless atom-optical field coupling strength $g$ together. Besides, we find that there are phase driven first-order and second-order superradiance phase transitions, and the quantum criticality for the second-order superradiance phase transition is studied. In addition, the experimental feasibility is discussed. This work will stimulate the studies of non-Hermitian superradiance quantum phase transitions and their experimental realizations, as well as the underlying universality class of phase transitions.
pacs:
64.60.-i, 05.70.Fh, 64.60.Fr
I Introduction
The quantum Rabi model (QRM) named as Rabi Rabi ; book describes the interaction of a two-level atom and a single-model optical field and can be viewed as the minimal version of the Dicke model Dicke . In the past years, there are efforts in finding the exactly analytically solutions of the QRM am_1 ; am_3 ; am_2 ; am_4 ; am_5 ; am_6 ; am_7 ; am_8 . Employing the Bargmann space representation bs , Braak obtained the transcendental function, namely the G-function, and found that the poles of the G-function were the exact solutions of QRM es1 . Quickly, Chen et al. obtained the exact solutions of a two-photon QRM by combining the Bogoliubov transformation and the G-function es2 . In the decade, the two methods have been used for calculating the exact solutions of the generalized QRMs ext_1 ; ext_2 ; ext_3 ; ext_4 ; ext_5 ; ext_6 ; ext_7 ; ext_8 ; ext_9 ; ext_10 ; ext_11 . It was found that the symmetry breaking will lead to the topological transition in the QRM tp1 and the generalized QRM tp2 .
Accompanying by the search for the exact solutions, there are interests in studying the quantum criticality of the QRMs or the generalized QRMs HP ; ps1 ; ps2 ; R_d1 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 ; zhangyy_2 . Accompanied by the quantum phase transition, i.e., the superradiance phase transition (SPT), there will occur macroscopic photon excitations R_d1 . In 2015, Hwang et al. studied the universal dynamics and the critical behaviors of the QRM and the exact phase boundary of SPT was obtained by an effective low-energy approximation HP . They found that the critical exponent of SPT $\nu$ is $\nu=\frac{3}{2}$ HP . By analyzing a general scaling function, Liu et al. found that the $\nu$ in the anisotropic QRM sill satisfies $\nu=\frac{3}{2}$ and the phase boundary of SPT is obtained by a semi-class method ps2 . Quickly, Wei et al. numerically verified the exact $\nu$ by analyzing the critical behaviors of the finite frequency fidelity susceptibility qc_1 . Subsequent researches show that for a multi-cavity system with two qc_2 ; qc_3 or three zhangyy_2 cascaded QRMs, the crtical exponent is still equal to that of a single Rabi model, but for a three-level QRM, there occurs new critical exponent with $\nu=1$zhangyy_3 . We notice that the SPTs in the QRMs more or less depend on the atom-optical field coupling strength. It motivate us whether there exist an unconventional SPT which is independent of the atom-optical field coupling.
Recently, the combination of the non-Hermiticity and the QRMs triggered intriguing physics. Cao et al. constructed a generalized non-Hermitian integrable QRM. They obtained the exact energy spectrum and eigenstates of this model by employing the Bethe ansatz es_NH_1 . The work provides a generalized strategy to engineer integrable non-Hermitian spin-boson models. By introducing the parity-time ($\mathcal{PT}$) symmetric gain and loss in the semi-classical Rabi model, Lee et al. PT_Rabi_1 found that there was $\mathcal{PT}$ symmetry breaking PT ; PT1 accompanied by the real-complex transition in energy. With the exact solutions of the $PT$-symmetric Rabi models, one can locate the $PT$-symmetry breaking point, i.e., the exceptional point (EP) EP . Xie et al. investigated the Floquet $\mathcal{PT}$-symmetric semi-classical Rabi model es_NH_2 . Under the multiple-photon resonance condition, they obtained the positions of the EPs from the exact Floquet spectrum. Employing the adiabatic approximation, Lu et al. obtained the exact energy spectrum of the $\mathcal{PT}$-symmetric QRM, by which the EPs were located es_NH_3 . With the emergence of EP, observables present discontinuous behaviors. By the Bogoliubov transformation and G-function, Li et al. solved the exact energy spectra for a one-photon and two-photon $\mathcal{PT}$-symmetric QRM, and the zeros of G-function located the EPs es_NH_4 . These works provide an inspiration for us to study whether there will be a completely real energy spectrum in a generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In addition, we will study the quantum criticality of the model and discuss the similarity and difference with Hermitian cases.
The paper is organized as follows. In Sec. II we introduce the generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In Sec. III, we obtain the exact phase boundary of the model. In Sec. IV, we study the observables which signal the SPT. In Sec. V, we discuss the quantum criticality of the SPT. A experimental feasibility is discussed in Sec. VI, and a summary is presented in Sec. VII.
II Two coupled non-Hermtian Rabi cavities
The generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry is constructed by two cascaded non-Hermitian Rabi cavities, whose Hamiltonian ($\hbar=1$) is given as
$H=\sum_{j=1,2}h_{j}+J(a^{{\dagger}}_{1}+a_{1})(a^{{\dagger}}_{2}+a_{2}),$ (1)
where $h_{j}$ describes non-Hermitian Rabi cavity $h_{j}=\omega_{p}a^{{\dagger}}_{j}a_{j}+\frac{\omega_{a}}{2}\sigma^{j}_{z}-\lambda e^{i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}$ with $j$ the cavity index and $a^{{\dagger}}_{j}$ ($a_{j}$) the creation (annihilation) of photons. $\omega_{p}$ is the optical field frequency, $\omega_{a}$ is the atomic frequency, and $\lambda e^{i\varphi}$ is the coupling strength between the atom and optical field with $\lambda$ the weight of strength and $\varphi$ the coupling phase. The non-Hermiticity of the system requires $\varphi\in(-\pi,0)\cup(0,\pi)$.
Figure 1: (Color online) Ground-state phase diagram in the $\mathcal{J}$-$g$ parameter space. (a) $\varphi=\frac{\pi}{2}$. (b) $\varphi=\frac{\pi}{4}$. The blue solid lines denote the phase boundary, satisfying $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$.
For a single non-Hermitian Rabi cavity, we argue that there is no parity-time ($\mathcal{PT}$) symmetry unless the coupling phase $\varphi=\pm\frac{\pi}{2}$. The time-reversal operator $\mathcal{T}$ dominates the complex conjugate operation, resulting in $\mathcal{T}a^{{\dagger}}_{j}(a_{j})\mathcal{T}$=$a^{{\dagger}}_{j}(a_{j})$. The parity operator $\mathcal{P}$ dominates the inversion operation, i.e., $\mathcal{P}\hat{x}_{j}\mathcal{P}$= $-\hat{x}_{j}$, and $\mathcal{P}\hat{p}_{j}\mathcal{P}$= $-\hat{p}_{j}$ PT . Thus, we have
$\mathcal{PT}h_{j}\mathcal{TP}$ $=\omega_{p}a^{{\dagger}}_{j}a_{j}-\mathcal{P}\left[\lambda e^{-i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}\right]\mathcal{P}+\frac{\omega_{a}}{2}\sigma^{j}_{z}$ (2)
$=\omega_{p}a^{{\dagger}}_{j}a_{j}+\lambda e^{-i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}+\frac{\omega_{a}}{2}\sigma^{j}_{z}.$
Only if $\varphi=\pm\frac{\pi}{2}$, the single non-Hermitian Rabi cavity is $\mathcal{PT}$-symmetric (see another prove in Ref. es_NH_4 ). For other values, it is not $\mathcal{PT}$-symmetric.
For the cascaded non-Hermitian QRM, we demonstrate that this system is completely not $\mathcal{PT}$-symmetric. $\mathcal{P}$ dominates a mirror operation between the two cavities, i.e., $\mathcal{P}a^{{\dagger}}_{j}(a_{j})\mathcal{P}=a^{{\dagger}}_{j^{\prime}}(a_{j^{\prime}})$, $\mathcal{P}\sigma^{j}_{z}(\sigma^{j}_{x})\mathcal{P}=\sigma^{j^{\prime}}_{z}(\sigma^{j^{\prime}}_{x})$ (here $j\neq j^{\prime}$), and $\mathcal{T}$ leads to $\mathcal{T}i\mathcal{T}=-i$. Performing $\mathcal{PT}$ on the cavity coupling term, we have
$\mathcal{PT}J\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right)\mathcal{TP}=J\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right).$ (3)
Performing $\mathcal{PT}$ on $h_{j}$, we have
$\mathcal{PT}h_{j}\mathcal{TP}$ $=\omega_{p}a^{{\dagger}}_{j^{\prime}}a_{j^{\prime}}-\lambda e^{-i\varphi}\left(a_{j^{\prime}}+a^{{\dagger}}_{j^{\prime}}\right)\sigma^{j^{\prime}}_{x}+\frac{\omega_{a}}{2}\sigma^{j^{\prime}}_{z}$ (4)
$\neq h_{j^{\prime}}.$
Therefore, here the non-Hermitian cascaded QRM is not $\mathcal{PT}$-symmetric.
III Exact and unconventional superradiance phase transition
To analytically solve the phase boundary of the ground-state superradiance phase transition, we introduce the dimensionless parameters $\eta=\frac{\omega_{a}}{\omega_{p}}$, $g=\frac{2\lambda}{\sqrt{\omega_{a}\omega_{p}}}$, and $\mathcal{J}=\frac{J}{\omega_{p}}$. Then, the Hamiltonian in Eq. (1) can be rewritten as
$\frac{H}{\omega_{p}}=\sum_{j=1,2}\frac{h_{j}}{\omega_{p}}+\mathcal{J}\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right),$ (5)
in which
$\frac{h_{j}}{\omega_{p}}=a^{{\dagger}}_{j}a_{j}+\frac{\eta}{2}\sigma^{j}_{z}-\frac{g\sqrt{\eta}}{2}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}.$ (6)
Performing the following transformation
$x_{j}$ $=\left(a^{{\dagger}}_{j}+a_{j}\right)/\sqrt{2\eta},$ (7)
$y_{j}$ $=i\left(a^{{\dagger}}_{j}-a_{j}\right)\sqrt{\eta/2},$
and dividing both sides of Eq. (5) by $\eta$, we obtain
$\mathcal{H}=\frac{H}{\omega_{p}\eta}\equiv\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}+\frac{y^{2}_{j}}{\eta^{2}}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right)+2\mathcal{J}x_{1}x_{2}.$ (8)
As studied in Refs. ps1 ; ps2 , the SPT appears in the under the infinite $\eta$ limit. Under this condition, we can obtain the analytical phase boundary of SPT. When $\eta\rightarrow\infty$, $\mathcal{H}$ reduces to
$\mathcal{H}=\frac{1}{2}\sum_{j=1,2}\left[x^{2}_{j}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right]+2\mathcal{J}x_{1}x_{2}.$ (9)
By diagonalizing $\mathcal{H}$, we obtain the energies $E_{\pm}(x_{1},x_{2})$ as
$E_{\pm}(x_{1},x_{2})=\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}\pm\sqrt{1+2g^{2}x^{2}_{j}}\right)+2\mathcal{J}x_{1}x_{2}.$ (10)
Performing the Taylor expansion on $E_{\pm}(x_{1},x_{2})$, and truncating them to the quadratic term $\mathcal{O}(x^{4}_{j})$, the low-energy branch $E_{-}(x_{1},x_{2})$ (contains the ground-state information) becomes
$E_{-}(x_{1},x_{2})=\frac{1}{2}\psi^{T}\Lambda\psi,$ (11)
in which $\psi=(x_{1},x_{2})^{T}$ and
$\Lambda=\begin{pmatrix}1-g^{2}e^{i2\varphi}&2\mathcal{J}\\ 2\mathcal{J}&1-g^{2}e^{i2\varphi}\end{pmatrix}.$ (12)
Evidently, the matrix $\Lambda$ consists of the real part ${\rm Re}(\Lambda)$ and the pure imaginary part ${\rm Im}(\Lambda)$ (which can be considered as the background loss), i.e., $\Lambda={\rm Re}(\Lambda)+{\rm Im}(\Lambda)$, with
${\rm Re}(\Lambda)$ $=\left[1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)\right]\sigma_{0}+2\mathcal{J}\sigma_{x},$ (13)
${\rm Im}(\Lambda)$ $=-i2g^{2}\cos\varphi\sin\varphi\sigma_{0}.$
Figure 2: (Color online) The real part of the ground-state energy ${\rm Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ under different $\eta$ with $g=0.7$ and $\phi=\frac{\pi}{2}$ in (a) and with $g=0.4$ and $\varphi=\frac{\pi}{4}$ in (b). (c) The second-order derivative of ${\rm Re}(E_{0})/\eta$ for $g=0.7$ and $\phi=\frac{\pi}{2}$. (d) The first-order derivative of ${\rm Re}(E_{0})/\eta$ for $g=0.4$ and $\phi=\frac{\pi}{4}$. The photonic truncation number is $N=80$.
In fact, the phase boundary of SPT is nothing to do with background loss ${\rm Im}(\Lambda)$, but is determined by ${\rm Re}(\Lambda)$. After diagonalizing ${\rm Re}(\Lambda)$, the eigenvalues ϵ±subscriptitalic-$\epsilon_{\pm}$ of ${\rm Re}(\Lambda)$ is obtained as $\epsilon_{\pm}=1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)\pm 2\mathcal{J}$. When the low-energy branch in ϵ±subscriptitalic-$\epsilon_{\pm}$ equals to zero, i.e, $\epsilon_{-}=0$, we can extract the phase boundary of the superradiance phase transition, and it satisfies $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$. From the expression of the phase boundary, we can see that there are special phases, i.e., $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, at which the phase boundary is fixed at $\mathcal{J}=\frac{1}{2}$, and is independent of the atom-optical field coupling parameter $g$. Such type of SPT independent of atom-optical field coupling is counter-intuitive and has not been reported before, so we name this type of SPT as unconventional SPT. While $\gamma$ is taken at other values, the phase boundary depends on the $g$ and $\mathcal{J}$ together. Thus, we name this atom-optical field coupling dependent SPT as the conventional SPT. Without loss of generality, in the following, we will study the energy, photon population, qubit population, the fidelity and the fidelity susceptibility in the ground state at $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ to characterize the properties of the SPT of the non-Hermitian two-Rabi cavity system. The ground state here refers to the state with the smallest real part of the energy. Intuitively, the phase diagrams for $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ are plotted in Fig. 1(a) and 1(b), respectively. The red regions denote the superradiance phases, separated from the normal phases (yellow regions) by the phase boundaries (blues curves).
Figure 3: (Color online) The ground-state photon population $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ as a function of $\mathcal{J}$. (a) $g=0.5$ and $\phi=\frac{\pi}{2}$; (b) $g=0.7$ and $\varphi=\frac{\pi}{2}$; (c) $g=0.4$ and $\varphi=\frac{\pi}{4}$; (d) $g=0.6$ and $\varphi=\frac{\pi}{4}$. The photonic truncation number is $N=80$ and $\eta=1500$.
IV Signatures of the superradiance phase transition
We detect the signatures and check the phase boundary of the SPT. We start by studying the ground-state energy to reveal the SPT. The energies are sorted by the real parts of the energies, and the ground-state energy means the energy with the lowest real part. Under $\varphi=\frac{\pi}{2}$ and different $\eta$, the real parts of the ground-state energies ${\rm Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ for $g=0.7$ are plotted in Fig. 2(a). The photonic truncation number is taken at $N=80$. As can be seen that with the emergence of SPT, ${\rm Re}(E_{0})/\eta$ for different $\eta$ decrease from $-1$ to lower values. According to the obtained analytical result, the transition point under $\varphi=\frac{\pi}{2}$ and $g=0.7$ should be at $\mathcal{J}=0.745$. Intuitively, the SPT occurs at the transition points predicted by the exactly analytical result (shown with the vertical black line). Meanwhile, we calculate ${\rm Re}(E_{0})/\eta$ in the $\varphi=\frac{\pi}{4}$ case and ${\rm Re}(E_{0})/\eta$-$\mathcal{J}$ relations for $g=0.4$ are plotted in Fig. 2(b). Similar to the $\varphi=\frac{\pi}{2}$, when $\mathcal{J}$ exceeds to the transition point, ${\rm Re}(E_{0})/\eta$ for different $\eta$ in the $\varphi=\frac{\pi}{4}$ case decrease from $-1$ to lower values, signaling the appearance of SPT as well. From the analytical expression of phase boundary, it is known that the transition point under $\varphi=\frac{\pi}{2}$ should be at $J=0.5$ regardless of what value of $g$ is. It is readily seen that for $g=0.4$, the transition points of ${\rm Re}(E_{0})/\eta$ occur at the analytically obtained transition points. Although the ground state energy of the system decreases after the occurrence of conventional and unconventional SPTs, it should be pointed out that they do not belong to the same phase transition type. As shown in Fig. 1(c), the second-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is not continuous at the the transition point. Figure 1(d) shows that the first-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{4}$ and $g=0.4$. The results show that the conventional SPT should be the second-order quantum phase transition, while the unconventional SPT should be the first-order quantum phase transition. In fact, only the $|\varphi|=\frac{\pi}{4}$ and $|\varphi|=\frac{3\pi}{4}$ cases correspond to the second-order quantum phase transition, while the cases with other $\varphi$ correspond to the first-order quantum phase transition. However, we are more concerned with conventional and unconventional superradiance phase transitions in this paper, so the results of other $\varphi$ are not considered
Figure 4: (Color online) The ground-state atom population $\sigma^{1(2)}_{z}$ as a function of $\mathcal{J}$. (a) $g=0.5$ and $\phi=\frac{\pi}{2}$; (b) $g=0.7$ and $\varphi=\frac{\pi}{2}$; (c) $g=0.4$ and $\varphi=\frac{\pi}{4}$; (d) $g=0.6$ and $\varphi=\frac{\pi}{4}$. The photonic truncation number is $N=80$ and $\eta=1500$.
Except for the ground-state energy, the phenomenon of SPT can be readily seen from the ground-state photon population. Under $\varphi=\frac{\pi}{2}$, ground-state photon populations $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ for $g=0.5$ and $g=0.7$ are plotted in Fig. 3(a) and 3(b), respectively. It is seen that there is no photon excitation when $\mathcal{J}$ is smaller than the transition point, but there are macroscopic photon excitations when the cavity coupling parameter $\mathcal{J}$ exceeds the transition point. The phenomenon is similar to the $\varphi=\frac{\pi}{4}$ case. For $g=0.4$ and $g=0.6$, the corresponding ground-state photon populations are plotted in Fig. 3(c) and 3(d), respectively. We can see that there are macroscopic photon excitations when $\mathcal{J}$ is larger than $0.5$, while the photon excitations are evidently suppressed when $J<0.5$.
A reasonable explaination for the macroscopic photon excitations in the suprradiance phase is that the atoms are excited and become active. To check this point, the atom populations $\langle\sigma^{1(2)}_{z}\rangle$ as a function of $\mathcal{J}$ for different $\varphi$ and $g$ are plotted in Fig. 4(a)-4(d). We can see that when $\mathcal{J}$ is smaller than the transition point, the atoms are stable in the ground state, with populations $\langle\sigma^{1(2)}_{z}\rangle=-1$. When $\mathcal{J}$ crosses the transition point, it is seen that the populations of atoms deviate from -1, implying that the atoms are excited and become active. The excited atoms dominate the macroscopic photon excitations in the SPT. Although the photon excitations and the atom populations of the system vary obviously with occurrence of SPT, the two types of SPTs lead to different varying behaviors of observables. From Fig. 3, we can see that when the conventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ continuously varies to macroscopic value, while the unconventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ directly jumps to macroscopic values without asymptotic process. Similarly, we can see the obvious difference between the two types of SPTs from the atomic populations in Fig. 4.
Figure 5: (Color online) The excitation gaps ${\rm Re}(\Delta_{E})$ and ${\rm Im}(\Delta_{E})$ as functions of $\mathcal{J}$ under different $\eta$. (a1) and (a2): $g=0.5$ and $\phi=\frac{\pi}{2}$; (b1) and (b2): $g=1$ and $\varphi=\frac{\pi}{2}$; (c1) and (c2): $g=0.4$ and $\varphi=\frac{\pi}{4}$; (d1) and (d2): $g=0.6$ and $\varphi=\frac{\pi}{4}$. The photonic truncation number is $N=36$.
In addition to the photon population and the atom population, the excitation gaps behind the two types of SPTs are different as well. The excitation gaps here contain two meanings. One is the real excitation gap ${\rm Re}(\Delta_{E})/\eta$ between the real parts of the lowest two energies. Another is the imaginary excitation gap ${\rm Im}(\Delta_{E})/\eta$ between the imaginary parts of the lowest two energies. Figures. 5(a1)-5(d1) (Figs. 5(a2)-5(d2)) presents ${\rm Re}(\Delta_{E})/\eta$ (${\rm Im}(\Delta_{E})/\eta$) as a function of $\mathcal{J}$ under different $\eta$. We can see that no matter $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$, the real excitation gap experiences a gap closing process when $\mathcal{J}$ exceeds the transition point. However, for the imaginary excitation gap, there is an obvious difference between the two studied cases. From Figs. 5(a2) and 5(b2), we can see that ${\rm Im}(\Delta_{E})/\eta$ always keeps zero (${\rm Im}(\Delta_{E})/\eta$ far less than $10^{-14}$ can be numerically regarded as zero.) before and after the transition point. From Figs. 5(c2) and 5(d2), we can see that the imaginary excitation gaps experience the gap closing process, similar to the circumstances for the real excitation gaps.
Figure 6: (Color online) The maximal imaginary part of the energy spectrum ${\rm Im}(E)_{max}/\eta$ as functions of $\mathcal{J}$ under different $\eta$. (a) $\varphi=\frac{\pi}{2}$ and $g=0.7$. (b) $\varphi=\frac{\pi}{2}$ and $g=0.8$. (c) $\varphi=\frac{\pi}{4}$ and $g=0.4$. (d) $\varphi=\frac{\pi}{4}$ and $g=0.6$. The photonic truncation number is $N=36$.
It is noted that for a $\mathcal{PT}$-symmetric Rabi system, there exists real-complex transition in energy spectrum PT ; PT1 ; es_NH_2 ; es_NH_3 ; es_NH_4 . Surprisingly, for the non-Hermitian two-Rabi cavity system without $\mathcal{PT}$ symmetry, there exists real energy spectrum as well. Under $\varphi=\frac{\pi}{2}$, the maximal imaginary of the energy spectra for $g=0.7$ and $g=0.8$, i.e., ${\rm Im}(E)_{max}$ are plotted in Figs. 6(a) and 6(b), respectively. The analytically obtained superradiance transition point for $g=0.7$ is $J=0.745$, and that for $g=0.8$ is $J=0.82$. ${\rm Im}(E)_{max}/\eta$ approaches to zero before and after the superradiance transition point implying that the energy spectra under $\varphi=\frac{\pi}{2}$ are real, and there is no real-complex transition in energy spectrum. Compared to the $\varphi=\frac{\pi}{2}$ case, the spectral property of the $\varphi=\frac{\pi}{4}$ is different. Under $\varphi=\frac{\pi}{4}$, ${\rm Im}(E)_{max}/\eta$ for $g=0.4$ and $g=0.6$ are plotted in Figs. 6(c) and 6(d), respectively. Different from the $\varphi=\frac{\pi}{2}$ cases, here the energy spectra are complex because of the nonvanishing ${\rm Im}(E)_{max}/\eta$. In fact, the real and complex characteristics of the energy spectrum can be understood analytically as well. From Eq. (10) and Eq. (13), we know that when $\varphi=\frac{\pi}{2}$, the imaginary part disappears, but when $\varphi=\frac{\pi}{4}$, the imaginary part still exists. The results in Fig. 6 provide numerical validation of the previous analytical derivation. Considering the fact that the $\varphi=\frac{\pi}{2}$ cases have full real energy spectra, we conjecture that here the conventional SPT belongs to the same universality class as the Hermitian Rabi systems HP ; ps2 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 and the critical exponent of the phase transition can be obtained from the critical behavior. However, for the complex energy spectrum case ($\varphi=\frac{\pi}{4}$), we conjecture that there is no critical behavior. Although the existence of quantum criticality of conventional SPT and the absence of quantum criticality of unconventional SPT have been reflected in the photon population and atom population, we still want to further verify them by studying the ground-state fidelity and fidelity susceptibility.
Figure 7: (Color online) Fidelity susceptibility $\chi_{\mathcal{F}}$ as functions of $\mathcal{J}$ under different $\eta$ with $\varphi=\frac{\pi}{2}$ and $g=0.7$ in (a), and with $\varphi=\frac{\pi}{2}$ and $g=0.8$ in (b). The involved parameter is $\delta\mathcal{J}=10^{-5}$. Fidelity $\mathcal{F}$ as functions of $\mathcal{J}$ under different $\eta$ with $\varphi=\frac{\pi}{4}$ and $g=0.4$ in (c), and with $\varphi=\frac{\pi}{4}$ and $g=0.6$ in (c). The involved parameter is $\delta\mathcal{J}=5\times 10^{-4}$. The photonic truncation number in all calculations is $N=80$.
V conventional and unconventional superradiance phase transition
We focus on the quantum phase transition and its critical behavior triggered by the cavity coupling parameter $\mathcal{J}$. The fidelity for the ground state $\mathcal{F}$ is defined as fidelity_0 ; fidelity_1 ; fidelity_2 ; fidelity_3
$\mathcal{F}=|\langle\psi_{0}(\mathcal{J})|\psi_{0}(\mathcal{J}+\delta\mathcal{J})|,$ (14)
which weighs the survival probability of the ground state after the cavity coupling parameter $\mathcal{J}$ from an infinitesimal distance $\delta\mathcal{J}$. Further, according to Refs. FS_1 ; FS_2 ; FS_3 ; FS_4 ; FS_5 , we can obtain the fidelity susceptibility $\chi_{\mathcal{F}}$ as
$\chi_{\mathcal{F}}=\lim_{\delta\mathcal{J}\rightarrow 0}\frac{-2\ln\mathcal{F}(\mathcal{J},\delta\mathcal{J})}{(\delta\mathcal{J})^{2}}.$ (15)
If there exists quantum phase transition, the fidelity will experience a drop at the phase transition point, and the fidelity susceptibility will peak at the phase transition point.
Figure 8: (Color online) (a) Universal finite $\eta$ scaling relation of the maximal fidelity susceptibility $\chi^{max}_{F}$. (b) Data collapse of $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}$. The involved parameters are $\varphi=\frac{\pi}{2},$$g=0.7$, and $\delta J=10^{-5}$.
Under $\varphi=\frac{\pi}{2}$, $\chi_{\mathcal{F}}/\eta$ for $g=0.7$ and $g=0.8$ are plotted in Figs. 7(a) and 7(b), respectively. As seen that with the increase of $\mathcal{J}$, $\chi_{\mathcal{F}}/\eta$ smoothly reaches the peak and then smoothly decreases. Besides, $\chi_{\mathcal{F}}/\eta$ presents obvious critical behaviors, because it is seen that with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ at the coupling parameter $\mathcal{J}_{max}$ gradually tends to the exact solution. Moreover, with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ get higher. However, for the $\varphi=\frac{\pi}{4}$ cases, there is no critical phenomenon. Under $\varphi=\frac{\pi}{4}$, the fidelity $\mathcal{F}$ for $g=0.4$ and $g=0.6$ are plotted in Figs. 7(c) and 7(d), respectively. Intuitively, $\mathcal{F}$ sharply drop to zero at single parameter $\mathcal{J}$ and there is no asymptotic behavior between the valley of $\mathcal{F}$ (or the peak of $\chi_{\mathcal{F}}$) and $\eta$. In other words, the conventional SPT exists quantum criticality while there is no quantum criticality for the unconventional SPT.
Next, we study the quantum criticality of the conventional SPT ($\varphi=\frac{\pi}{2}$ case). From the $\chi_{\mathcal{F}}$-$\eta$ curves, the maximal fidelity susceptibility $\chi^{max}_{F}$ under a specific $\eta$ is obtained. Then, the universal finite $\eta$ scaling relation of $\chi_{\mathcal{F}}$ for $g=0.7$ is plotted in Fig. 8(a). The black fitting curve satisfies $\ln\chi_{\mathcal{F}}=(-1.363\pm 0.004)\ln\eta+2.388$. It implies the scaling exponent $\mu$ of $\chi_{\mathcal{F}}$ is $\mu=-1.363\pm 0.004$. As introduced in Ref. qc_1 ; RMP2011 , the relationship between $\mu$ and $\nu$ obeys $\mu=\frac{2}{\nu}$, by which we obtain $\nu=1.4674\pm 0.004$. Although the numerically obtained critical exponents $\mu$ and $\nu$ has slightly numerical deviation from the exact solution (with $\nu_{exact}=3/2$) of the single Rabi model HP ; ps2 , we argue that the quantum phase transition in the $\varphi=\frac{\pi}{2}$ is the same universality class as that in the single Rabi model. Within the neighborhood of the phase transition point, $\chi_{\mathcal{F}}$ scales as $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}=f\left[\eta^{1/\nu}\left(\mathcal{J}-\mathcal{J}_{max}\right)\right]$, where $f(x)$ is a universal scaling function. Under different $\eta$, we find the corresponding data of $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}$ are collapsed into one curve with the exact critical exponent $\nu=\frac{3}{2}$, signaling the same universal class as the single Rabi model HP ; ps2 . In fact, for the Dicke model Dicke_1 ; Dicke_2 and the Lipkin-Meshkov-Glick model LMG_1 ; LMG_2 ; LMG_3 , the critical exponent satisfies $3/2$ as well, therefore the conventional SPT of the non-Hermitian cascade Rabi cavities we have studied belongs to the same universality class as those in the above mentioned systems.
VI Experimental feasibility
The experimental exploration of the phenomena discussed here could be facilitated by a proposed circuit-QED architecture, as detailed in Ref. SC . This setup comprises two resonators, each equipped with an embedded artificial atom, specifically a superconducting flux qubit. The key feature of this design is the implementation of a coupling mechanism between the resonators that is both broadly and rapidly tunable. By leveraging fast-modulating fields, this tunable coupling can effectively emulate a phase-dependent interaction term, a crucial component for investigating PT-related physics. Furthermore, the architecture allows for the coupling between the flux qubits and the resonators to be pushed into the ultrastrong coupling regime ($g>0.1$), and even the deep-strong coupling regime ($g>1$). These advanced capabilities of the circuit-QED system provide a solid platform not only for testing the theoretical predictions presented in this article but also for realizing these concepts in practical, soon-to-be-available devices.
VII Summary
Herein, a non-Hermitian QRM consisting of two cascaded non-Hermitian Rabi cavities has been studied. Similar to Hermitian QRMs, the superradiance phase boundary of the non-Hermitian one can be exactly solved as well. Accompanied by the emergence of SPT, the ground state energy gets lower. Meanwhile, the atoms become active and radiating photons outward. However, the non-Hermiticity leads to the exotic phenomena without Hermitian counterpart. When the atom-optical field coupling phase satisfies $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, the superradiance phase boundary is surprisingly dominated by a constant independent of the strength of the atom-cavity field coupling, suggesting the unconventional SPT. For other coupling phases, the superradiance phase boundary depends on the cavity coupling strength and the atom-optica field coupling strength together, suggesting the conventional SPT. Meanwhile, we have argued that the conventional SPT belongs to the second-order quantum phase transition, while the unconventional SPT belongs to the first-order quantum phase transition. Even though the model we have studied is not $\mathcal{PT}$-symmetric, there still exists fully real energy spectrum. By the photon population, atom population, and the fidelity we have known that there is no quantum criticality for the unconventional SPT due to characteristics of the absence of asymptotic process and the single-parameter sharp drop of fidelity. Differently, for conventional SPT, there exists quantum criticality, and the critical exponent extracted from the scaling analyze of fidelity susceptibility show that the SPT belong to the same universality class as the single Rabi model. With the development of circuit-QED experimental technique, the found SPTs may be realized in high-performance circuit- QED architecture.
We acknowledge the support from NSFC under Grant No. 12174346.